OpenStudy (anonymous):

Find an exact value. tan 75°

OpenStudy (anonymous):

I got five questions wrong in my assignment and I kind of want to see where I went wrong. :) any help is very appreciated

OpenStudy (anonymous):

\(75=30+45\) so this is a set up for the "addition angle formula" for tangent

OpenStudy (anonymous):

I think it may be that I didn't simplify...I chose 2 + sqrt 3 It's wrong

OpenStudy (jdoe0001):

heheh

OpenStudy (anonymous):

hello again :) I called my teacher but she wouldn't pick up...so I thought I'd try again and see what I did wrong

OpenStudy (anonymous):

God I despise pre-calc...I miss algebra...

OpenStudy (jdoe0001):

this one, is EXACTLY more or less to the one you did at http://openstudy.com/study#/updates/51e994d1e4b0d999840eb6ac the one there uses \(\bf tan(45 - 30) = \cfrac{tan(45) - tan(30)}{1 + tan(45)tan(30)}\) this one uses \(\bf tan(45 + 30) = \cfrac{tan(45) + tan(30)}{1 - tan(45)tan(30)}\) same values you'll be adding and simplifying

OpenStudy (anonymous):

What I don't about the last one is how that person got the 2 out of nowehere...how do you get that number here?

OpenStudy (jdoe0001):

hehe, well, what part confused you on the simplifying?

OpenStudy (anonymous):

I'm not very good at fractions...so that part of it...you cross multiplied right?

OpenStudy (jdoe0001):

well, no that much, mainly adding and cancelling like-terms in the numerator and denominator